# GPS-IDS: An Anomaly-based GPS Spoofing Attack Detection Framework for Autonomous Vehicles

### v1

@@ Abstract
Autonomous Vehicles (AVs) heavily rely on sensors and communication networks like Global Positioning System (GPS) to navigate autonomously. Prior research has indicated that networks like GPS are vulnerable to cyber-attacks such as spoofing and jamming, thus posing serious risks like navigation errors and system failures. These threats are expected to intensify with the widespread deployment of AVs, making it crucial to detect and mitigate such attacks. This paper proposes GPS Intrusion Detection System, or GPS-IDS, an Anomaly Behavior Analysis (ABA)-based intrusion detection framework to detect GPS spoofing attacks on AVs. The framework uses a novel physics-based vehicle behavior model where a GPS navigation model is integrated into the conventional dynamic bicycle model for accurate AV behavior representation. Temporal features derived from this behavior model are analyzed using machine learning to detect normal and abnormal navigation behavior. The performance of the GPS-IDS framework is evaluated on the AV-GPS-Dataset — a real-world dataset collected by the team using an AV testbed. The dataset has been publicly released for the global research community. To the best of our knowledge, this dataset is the first of its kind and will serve as a useful resource to address such security challenges.
@@
Autonomous Vehicle, Anomaly Detection, AV-GPS-Dataset, GPS Attacks, Intrusion Detection System, Physics-based Behavior Modeling, Modified Dynamic Bicycle Model.
@@ I Introduction
Vehicles are getting increasingly autonomous and becoming ever more reliant on onboard sensors and communication networks, aiming to make transportation faster, safer, and environmentally sustainable by reducing human intervention in driving tasks [1, 70]. Researchers have shown human error to be the cause of over 90% accidents, resulting in 40 thousand deaths and over 2 million injuries annually in the United States alone [2, 3] — a statistic that will be reduced with the transition to AVs [4]. AVs rely on Automotive Sensing, a collection of diverse sensors that enable environmental perception and safe navigation without requiring constant human input [5]. Automotive sensing is broadly categorized into three types: Self-sensing, Surrounding-sensing, and Localization [5]. Self-sensing refers to how an autonomous vehicle gathers and interprets information about its own state, including its position, velocity, and acceleration. Surrounding-sensing is the ability of a vehicle to perceive its environment, like recognizing traffic signs, understanding weather conditions, or measuring the state of other vehicles around it. Localization determines the local and global positions of the vehicle and helps it to navigate safely to the desired destination. For self-sensing and surrounding sensing, AVs use an array of sensors like Inertial Measurement Units (IMUs), gyroscopes, odometers, Controller Area Network (CAN) bus, cameras, Light Detection and Ranging (LiDARs), etc. [6]. For localization and navigation, they rely on satellite-based navigation systems like Global Navigation Satellite System (GNSS) [7].
@@
Motivated by these challenges, this paper presents GPS-IDS, an Anomaly Behavior Analysis-based Intrusion Detection System that uses a novel physics-based vehicle behavior model — a modification of the conventional dynamic bicycle model, to detect GPS spoofing attacks on AVs. Temporal features extracted from this vehicle behavior model are used to capture the normal behavior of the AV. Afterward, Machine Learning models are employed to detect the modeled normal behavior from abnormal behavior.
@@
  • The paper presents the Autonomous Vehicle Behavior Model— a modified dynamic bicycle model that integrates an autonomous GPS navigation model. This modification accurately represents the normal navigation behavior of an AV, capturing the lateral dynamics and states. Temporal features extracted from the AV Behavior Model are used to observe the vehicle’s normal behavior. This normal behavior is separated from the attacks using machine learning techniques.
  • The paper introduces the “AV-GPS-Dataset” that captures 44 features of GPS-guided navigation of AVs with and without cyber-attacks. In contrast to the existing related datasets collected from simulated environments [8, 10, 11], this dataset is collected from practical field experiments with real-world GPS spoofing attacks performed on an autonomous vehicle testbed.
  • The proposed GPS-IDS framework is evaluated on the AV-GPS-Dataset, showing an F1 score of up to 94.4% with up to 13s detection time improvement compared to the Extended Kalman Filter (EKF) detector implemented in the experimental setup.
The rest of the paper is organized as follows: Section II discusses the related work; Section III introduces the GPS-IDS framework and explains the Autonomous Vehicle Behavior Model; Section IV presents the experimental evaluation of the GPS-IDS framework along with the dataset details, and Section V concludes the paper.
II Related Work
@@ II-A1 Physics-based Model
A physics-based vehicle model is a mathematical representation that simulates the vehicle’s physical behavior using factors like kinematics, dynamics, motion, forces, and environmental interactions. In [15], the authors adopted a physics-based dynamic vehicle model incorporating an EKF estimator to determine the vehicle’s longitudinal and lateral velocities and yaw rate. In [30], Kong et al. explored two physics-based vehicle models, namely kinematic and dynamic bicycle models, for model-based controller design in autonomous driving and presented a comparison between them using experimental data. Several research contributions have employed software tools such as Modelica, Robot Operating System (ROS), Gazebo, and Simulink to simulate physics-based vehicle models or their components. Notable applications include racing car modeling [16], vehicle drivability modeling [9], heterogeneous physical system modeling, and simulating AV testbeds [17].
@@ II-B Intrusion Detection System (IDS)
Intrusion Detection Systems or IDSs are software designed to detect cyber-attacks and can be classified into four types [12]: 1) Signature-based IDS, 2) Anomaly-based IDS, 3) Specification-based IDS, and 4) Hybrid IDS. Signature-based IDSs rely on a pre-learned attack signature database, where each observation matched with an entry in the attack database is considered as an attack. Signature-based IDSs heavily rely on an up-to-date attack signature database and are unable to detect new or modified attacks [13]. An Anomaly-based IDS relies on modeling techniques to capture the system’s normal behavior, wherein any observed behavior outside the model-defined norm is classified as malicious. This reliance on normal behavior models allows Anomaly-based IDSs to detect new or modified attacks at the cost of increased modeling complexity [14]. Specification-based IDSs use a set of rules and policies that define the expected behavior of different system components such as sensor nodes or motor commands. Hybrid IDSs are a combination of Signature-based, Anomaly-based, and Specification-based IDSs. The Anomaly-based IDS approach presented in this work detects GPS spoofing attacks on AVs by utilizing an extensive physics-based AV behavior model and real-life vehicular data, aiming to use a more accurate operational baseline and overcome the high error rates present in Anomaly-based IDSs.
@@ II-D GNSS Security
In this section, we highlight the research aimed at detecting GNSS attacks. GNSS security literature can be classified into two broad categories: (1) GPS signal characteristics-based approach [21], [22], [23], [24], [25], [26], and (2) Machine learning-based approach [65], [69] [66], [10], [67], [68]. The GPS signal characteristics-based approaches rely on signal processing techniques to detect attacks. For instance, in [21], Psiaki and Humphreys proposed a GPS spoofing attack detection scheme based on the direction-of-arrival (DOA) induction. He et al. [23] proposed the use of GPS signal distortion to detect GPS spoofing attacks. For connected and autonomous vehicles, an anomaly detection model is proposed by Yang et al. [24] to detect GPS spoofing on the localization system using the “Learning From Demonstration” technique. Milaat and Liu [25] proposed a decentralized technique where vehicles exchange GPS code pseudo-range values with neighbors using short-range communications to detect high correlations during spoofed GPS signal arrival. On the other hand, machine learning-based approaches rely on machine learning and data analytics techniques to detect attacks. Researchers have used one-class classifiers [10], Artificial Neural Networks [65], Long Short Term Memory (LSTM) networks [66], [67] to detect GPS attacks. However, the majority of these works adopt a segregated approach focused on sensor node-specific detection techniques and either use a small (and restrictive) GPS dataset, or use an attack dataset collected from simulated environments using tools like Matlab and Gazebo, lacking the representation of the real world.
This paper overcomes these limitations by introducing an anomaly-based intrusion detection approach and using real-world data. The approach relies on a novel modified dynamic bicycle model with GPS navigation to model the normal behavior of an AV. Temporal features highlighted in this behavior model are extracted from real-world experiments to model the vehicle’s normal navigation behavior. Machine learning techniques are then used to classify the normal navigation behavior from the abnormal.
@@ III GPS Intrusion Detection System (GPS-IDS)
GPS-IDS is an Anomaly Behavior Analysis-based Intrusion Detection System designed to identify GPS attacks on AVs by continuously monitoring the behavior of the vehicle and considering any deviation from the expected behavior as anomalous. The framework relies on a physics-based vehicle model, that is a modified dynamic bicycle model, called the Autonomous Vehicle Behavior Model to represent the vehicle’s normal behavior, as the physics model captures the vehicle’ps state at any instantaneous time $t$. Features highlighted in the physics model are used in machine learning analysis to detect anomalous behavior of the vehicle. Fig. 1 shows the architecture of the GPS-IDS framework.
III-A Components of GPS-IDS
The GPS-IDS framework has two modules: the Data Collection Module, and the Anomaly Behavior Analysis Module.
III-A1 Data Collection Module
@@
This module collects the raw data from the AV system and stores it in a shared database. Raw data includes GPS signals and an ongoing stream of vehicle dynamics data, which are obtained from the onboard state-measuring sensors.
III-A2 Anomaly Behavior Analysis Module
This module reads the raw data from the shared database and performs Anomaly Behavior Analysis to classify the data as normal or abnormal.
III-B Anomaly Behavior Analysis (ABA) for GPS-IDS
[…]
$D(s)=\begin{cases}Abnormal\hskip 28.45274ptif\quad f_{norm}(s,M)>\mathbb{T}\\ Normal\hskip 36.98866ptotherwise\end{cases}$
Detection takes place when the detector $D$ classifies a sample as abnormal, regardless of whether it is genuinely an anomaly or a regular sample that has been wrongly classified as one. The detection errors are defined over a test set $U_{t}^{R}$ which is a subset of $U^{R}$, $U_{t}^{R}\subseteq U^{R}$. The detector considers two kinds of errors: False Positives and False Negatives. A False Positive detection occurs when a normal sample $s\in N^{R}$ is detected as an abnormal event and is defined as $\varepsilon^{+}=\{s\in N^{R}|D(s)=abnormal\}$, while a False Negative detection occurs when the detector classifies an abnormal sample $s\in A^{R}$ as a normal event (undetected anomalies), that is $\varepsilon^{-}=\{s\in A^{R}|D(s)=normal\}$. The objective of GPS-IDS is to tune the predefined threshold $\mathbb{T}$ so that the overall error is minimized. Particularly, we prioritize the minimization of False Negative errors, as these undetected attacks pose greater risks compared to false alarms in the context of autonomous driving.
III-C Autonomous Vehicle Behavior Model
@@
In order to establish a comprehensive behavior model of AVs, it is essential to consider four key components: (1) Perception and Localization, (2) State Estimation, (3) Motion Planning, and (4) Control [28]. Perception and Localization rely on a combination of internal and external sensors, such as IMUs, Cameras, LiDARs, GPS, etc., to collect information about the surrounding environment. The remaining three components require a mathematical representation of the vehicle that encompasses its dynamics and motion. This study focuses specifically on GPS-guided localization in AVs, aiming to investigate the impact of a spoofing attack on the vehicle’s dynamics and motion. Therefore, predicting the dynamic state of the vehicle becomes crucial. To achieve this, a simple 2 Degrees of Freedom (2 DOF) lateral dynamics model is considered, and the bicycle model proposed by Rajamani et al. [29] is employed. Fig. 2 depicts the dynamic bicycle model of a vehicle in a 2-dimensional inertial frame.
@@
III-C1 Dynamic Bicycle Model
@@
III-C2 Perception/ Localization
The GPS enables the vehicle to determine its position and localize to the destination. Since this paper focuses on presenting a GPS intrusion detection system, only GPS-based localization is considered, and camera or vision sensor-based perception is not taken into account. To ensure safe GPS-guided localization for autonomous vehicles, it is necessary to continuously monitor various parameters such as GPS latitude and longitude, signal quality, GPS Dilution of Precision (DOP), and the number of satellites the vehicle is locked with. The normal behavior of the GPS signal for vehicular localization can be modeled by constantly monitoring the following parameters:
@@ III-C2 Perception/ Localization
III-C3 State Estimation
One of the well-established solutions for estimating the state of nonlinear systems is the Extended Kalman Filter or EKF. It integrates measurements from multiple sensors with a system model to estimate the state of the system with improved accuracy [53]. In the context of a GPS-guided autonomous vehicle, the EKF utilizes multiple sensor measurements in conjunction with the dynamic vehicle model to make accurate predictions about the vehicle’s state, including its position, orientation, velocity, and other relevant parameters. This estimation procedure takes into account the nonlinear relationship between the measurements and the position, while also considering measurement noise. As highlighted in the Related Work section, EKF state estimator is capable of detecting sensor anomalies and attacks against autonomous vehicles by monitoring the measurement deviations of multi-sensor readings [48, 55]. A similar EKF dynamic state estimator has been employed in our vehicle model and experimental testbed. In later sections, it will be shown that the GPS-IDS approach is able to detect a GPS spoofing attack faster than the employed EKF state estimation-based detection approach, which is supported by experimental validation.
@@ III-C3 State Estimation
$\textbf{x}_{k+1}=f_{cd}(\textbf{x}_{k},u_{k},W_{k})$
$\textbf{y}_{k+1}=g_{cd}(\textbf{x}_{k},E_{k})$
@@
$F=\left.\frac{\partial f_{cd}}{\partial\textbf{x}}\right|_{\hat{x}_{k-1},u_{k}}$
$G=\left.\frac{\partial g_{cd}}{\partial\textbf{x}}\right|_{\hat{x}_{k}}$
Thus, the nonlinear system can be transformed into a linear system with the following equations:
$\textbf{x}_{k+1}=F\textbf{x}_{k}+W_{k}$
$\textbf{y}_{k+1}=G\textbf{x}_{k}+E_{k}$
The update process is as follows:
$\textbf{P}_{\left.{k+1}\right|{k}}=F_{k}\textbf{P}_{\left.k\right|k}F^{T}_{k}+\textbf{Q}$
$\textbf{S}_{\left.{k+1}\right|{k}}=G_{k}\textbf{P}_{\left.{k+1}\right|k}G^{T}_{k}+\textbf{R}$
$\textbf{K}_{\left.{k+1}\right|{k}}=\frac{\textbf{P}_{\left.{k+1}\right|k}G^{T}_{k}}{\textbf{S}_{\left.{k+1}\right|k}}$
$\textbf{P}_{\left.{k+1}\right|{k+1}}=(\textbf{I}-\textbf{K}_{\left.{k+1}\right|k}G_{k})\textbf{P}_{\left.{k+1}\right|k}$
where Q is the covariance matrix of the prediction noise, R is the covariance matrix of the observation noise, and I is an identity matrix. Thus:
$\textbf{x}_{\left.{k+1}\right|k}=F\hat{\textbf{x}}_{\left.k\right|k}+W_{k}$
$\textbf{y}_{\left.{k+1}\right|k}=G\hat{\textbf{x}}_{\left.k\right|k}+E_{k}$
@@
$\hat{\textbf{x}}_{\left.{k+1}\right|{k+1}}=x_{\left.{k+1}\right|k}+\textbf{K}_{\left.{k+1}\right|k}(\textbf{y}_{k+1}-\textbf{y}_{\left.{k+1}\right|k})$
III-C4 Motion Planning
@@
From 18, we can say that when the vehicle is traveling along the desired path, $\psi(t)=\psi_{t}(t)$, and $e(t)=0$. When the vehicle deviates from the desired path, $\psi(t)$ differs from $\psi_{t}(t)$ and the $e(t)$ increases accordingly. In general, an increase in the difference between the target yaw angle and the current yaw angle leads to a corresponding amplification in the cross-track error.
III-C5 Controller
@@
Derived from the discussed software components, the architecture of the Autonomous Vehicle Behavior Model is depicted in Fig. 4.

### v2

@@ Abstract
Autonomous Vehicles (AVs) heavily rely on sensors and communication networks like Global Positioning System (GPS) to navigate autonomously. Prior research has indicated that networks like GPS are vulnerable to cyber-attacks such as spoofing and jamming, thus posing serious risks like navigation errors and system failures. These threats are expected to intensify with the widespread deployment of AVs, making it crucial to detect and mitigate such attacks. This paper proposes GPS Intrusion Detection System, or GPS-IDS, an Anomaly-based intrusion detection framework to detect GPS spoofing attacks on AVs. The framework uses a novel physics-based vehicle behavior model where a GPS navigation model is integrated into the conventional dynamic bicycle model for accurate AV behavior representation. Temporal features derived from this behavior model are analyzed using machine learning to detect normal and abnormal navigation behaviors. The performance of the GPS-IDS framework is evaluated on the AV-GPS-Dataset— a GPS security dataset for AVs comprising real-world data collected using an AV testbed, and simulated data representing urban traffic environments. To the best of our knowledge, this dataset is the first of its kind and has been publicly released for the global research community to address such security challenges.
@@
Autonomous Vehicles, Anomaly Detection, CARLA, GPS Spoofing, Intrusion Detection System, Physics-based Behavior Modeling, Robotics, Secure Navigation.
@@ I Introduction
Vehicles are getting increasingly autonomous and becoming ever more reliant on onboard sensors and communication networks, aiming to make transportation faster, safer, and environmentally sustainable by reducing human intervention in driving tasks [1, 75]. Researchers have shown human error to be the cause of over 90% accidents, resulting in 40 thousand deaths and over 2 million injuries annually in the United States alone [2, 3] — a statistic that will be reduced with the transition to AVs [4]. AVs rely on Automotive Sensing, a collection of diverse sensors that enable environmental perception and safe navigation without requiring constant human input [5]. Automotive sensing is broadly categorized into three types: Self-sensing, Surrounding-sensing, and Localization [5]. Self-sensing refers to how an autonomous vehicle gathers and interprets information about its own state, including its position, velocity, and acceleration. Surrounding-sensing is the ability of a vehicle to perceive its environment, like recognizing traffic signs, understanding weather conditions, or measuring the state of other vehicles around it. Localization integrates internal self-sensing and surrounding sensing data with external information from satellite-based navigation systems to determine the local and global positions of the vehicle. For self-sensing and surrounding sensing, AVs use an array of sensors like Inertial Measurement Units (IMUs), gyroscopes, odometers, Controller Area Network (CAN) bus, cameras, Light Detection and Ranging (LiDARs), etc. [6]. For localization and navigation, they rely on satellite-based navigation systems like Global Navigation Satellite System (GNSS) [7].
@@
Motivated by these challenges, this paper presents GPS-IDS, an Anomaly-based Intrusion Detection System that uses a novel physics-based vehicle behavior model (a modification of the conventional dynamic bicycle model) and machine learning to detect GPS spoofing attacks on AVs. The bicycle model is a simplified representation of a vehicle’s dynamics, typically used in the context of autonomous vehicle control systems. It approximates the vehicle as a two-wheel system, capturing the essential dynamics without the complexity of a full multi-wheel model. By extracting temporal features from this model, GPS-IDS captures the normal behavior of AVs. Machine learning models are then utilized to differentiate between normal and abnormal behaviors. The detection of GPS spoofing attacks is achieved by identifying deviations from the normal behavior pattern, indicating potential intrusions.
@@
  • The paper presents the Autonomous Vehicle Behavior Model— a hybrid vehicle model that integrates a modified physics-based dynamic bicycle model with GPS navigation. This modification allows for an accurate representation of the normal navigation behavior of AVs. Temporal features extracted from the physics-based model are utilized in the data-centric analysis to detect abnormal behaviors using machine learning.
  • The paper introduces the AV-GPS-Dataset, including:
    • Real-world Data: Collected from field experiments with real GPS spoofing attacks on an AV testbed, capturing 44 features of GPS-guided navigation.
    • Simulation-generated Data: Generated using the CARLA (Car Learning to Act) [65] simulator, representing various urban scenarios with diverse traffic conditions. This part of the dataset helps to understand GPS spoofing effects in urban environments.
  • In contrast to the existing related datasets collected solely from simulated environments [8, 10, 11], the AV-GPS-Dataset features both real-world and simulated data, providing normal and GPS spoofing attack scenarios.
  • The GPS-IDS framework is evaluated using the AV-GPS-Dataset. The framework achieves:
    • An F1 score of up to 94.4%, with a 56.5% improvement in detection time compared to the existing EKF-based GPS/INS detector in real-world data.
    • An F1 score of up to 97.1% in simulation-generated data with urban traffic scenarios, demonstrating high effectiveness in diverse environments.
The rest of the paper is organized as follows: Section II discusses the background and related work; Section III introduces the GPS-IDS framework and explains the components; Section IV explains the attacker model and defines the problem statement, Section V presents the experimental evaluation of the GPS-IDS framework along with the dataset details, and Section VI concludes the paper and outlines future directions.
II Background and Related Work
This section outlines the relevant technical background on vehicle modeling techniques, various types of Intrusion Detection Systems, existing works on cyber-attacks and GNSS security for vehicles, and our motivations to address the current limitations.
@@ II-A1 Physics-based Model
A physics-based vehicle model is a mathematical representation that simulates the vehicle’s physical behavior using factors like kinematics, dynamics, motion, forces, and environmental interactions. In [15], the authors adopted a physics-based dynamic vehicle model incorporating an EKF estimator to determine the vehicle’s longitudinal and lateral velocities and yaw rate. In [30], Kong et al. explored two physics-based vehicle models, namely kinematic and dynamic bicycle models, for model-based controller design in autonomous driving and presented a comparison using experimental data. Several research contributions have employed software tools such as Modelica, Robot Operating System (ROS), Gazebo, and Simulink to simulate physics-based vehicle models or their components. [16], [9], [17].
@@ II-B Intrusion Detection System (IDS)
Intrusion Detection Systems (IDS) are software designed to detect cyber-attacks and can be classified into four types [12]: 1) Signature-based IDS, 2) Anomaly-based IDS, 3) Specification-based IDS, and 4) Hybrid IDS. Signature-based IDSs rely on a pre-learned attack signature database, where each observation matched with an entry in the attack database is considered as an attack. Signature-based IDSs heavily rely on an up-to-date attack signature database and are unable to detect new or modified attacks [13]. An Anomaly-based IDS relies on modeling techniques to capture the system’s normal behavior, wherein any observed behavior outside the model-defined norm is classified as malicious. This reliance on normal behavior models allows Anomaly-based IDSs to detect new or modified attacks at the cost of increased modeling complexity [14]. Specification-based IDSs use a set of rules and policies that define the expected behavior of different system components such as sensor nodes or motor commands. Hybrid IDSs are a combination of Signature-based, Anomaly-based, and Specification-based IDSs. The Anomaly-based IDS approach presented in this work detects GPS spoofing attacks on AVs by utilizing an extensive physics-based AV behavior model and real-life vehicular data, aiming to use a more accurate operational baseline and overcome the high error rates present in Anomaly-based IDSs.
@@ II-D GNSS Security
In this section, we highlight the research aimed at detecting GNSS attacks. GNSS security literature can be classified into two broad categories: (1) GPS signal characteristics-based approach [21], [22], [23], [24], [25], [26], and (2) Machine learning-based approach [70], [74] [71], [10], [72], [73]. The GPS signal characteristics-based approaches rely on signal processing techniques to detect attacks. For instance, in [21], Psiaki and Humphreys proposed a GPS spoofing attack detection scheme based on the direction-of-arrival (DOA) induction. He et al. [23] proposed the use of GPS signal distortion to detect GPS spoofing attacks. For connected and autonomous vehicles, an anomaly detection model is proposed by Yang et al. [24] to detect GPS spoofing on the localization system using the “Learning From Demonstration” technique. Milaat and Liu [25] proposed a decentralized technique where vehicles exchange GPS code pseudo-range values with neighbors using short-range communications to detect high correlations during spoofed GPS signal arrival. On the other hand, machine learning-based approaches rely on machine learning and data analytics techniques to detect attacks. Researchers have used one-class classifiers [10], Artificial Neural Networks [70], Long Short Term Memory (LSTM) networks [71], [72] to detect GPS attacks.
The majority of these works adopt a segregated approach focused on sensor node-specific detection techniques and either use a small (and restrictive) GPS dataset or an attack dataset collected solely from simulated environments using tools like Matlab and Gazebo, lacking the representation of the real world. This paper overcomes these limitations by:
(1) Introducing an approach that is evaluated on both real-world data including practical GPS spoofing scenarios, as well as simulated data in diverse urban environment conditions, and (2) Employing a hybrid vehicle behavior model, a combination of physics-based and data-centric models, to accurately represent the behavior of AVs. The physics-based component is a modified dynamic bicycle model that includes GPS navigation, and the data-centric component extracts temporal features from this model to model the vehicle’s normal navigation behavior and identify the deviations using machine learning — thereby detecting GPS attacks.
Unlike the traditional GNSS security approaches, our approach benefits from diverse data and the combination of both physics-based and data-centric models, giving the IDS a comprehensive view and allowing faster detection of anomalies like GPS attacks.
@@ III GPS Intrusion Detection System (GPS-IDS)
GPS-IDS is an Anomaly-based Intrusion Detection System designed to identify GPS spoofing attacks on AVs by continuously monitoring the behavior of the vehicle and considering any deviation from the expected behavior as anomalous. The framework relies on a hybrid vehicle model called the Autonomous Vehicle Behavior Model to represent the vehicle’s behavior using both physics-based and data-centric components at any instantaneous time $t$. Features highlighted in the physics-based component are used in machine learning analysis to detect anomalous behavior of the vehicle. Fig. 1 shows the architecture of the GPS-IDS framework. The framework has two modules: the Data Collection Module, and the Anomaly Behavior Analysis Module.
III-A Data Collection Module
This module collects the raw data from the AV system and stores it in a shared database. It ensures that the necessary data is collected in real-time and is available for subsequent analysis. It has 3 main components: 1) Vehicle System, 2) Onboard Data Storage, and 3) Autonomous Vehicle Behavior Model.
III-A1 Vehicle System and GPS Receiver
The vehicle system includes a GPS receiver that is susceptible to spoofing attacks, where an attacker can manipulate the GPS receiver to provide false location data to the vehicle.
III-A2 Onboard Data Storage
@@
The onboard data storage system of the vehicle temporarily stores the raw data collected by the vehicle system. Raw data includes GPS signals and an ongoing stream of vehicle dynamics data, which are obtained from the onboard state-measuring sensors. This storage acts as a buffer that holds data before being analyzed by the subsequent modules.
III-A3 Autonomous Vehicle Behavior Model
@@
In order to establish a comprehensive behavior model of AVs, it is essential to consider four key components: (a) Localization or Navigation, (b) State Estimation, (c) Motion Planning, and (d) Control [28]— and the Autonomous Vehicle Behavior Model is comprised of these four components. Localization relies on a combination of internal and external sensors (like GPS, IMUs, Cameras, LiDARs, etc.) to collect information about the surrounding environment. The remaining three components require a mathematical representation of the vehicle that encompasses its dynamics. This study focuses specifically on GPS-guided localization in AVs, aiming to investigate the impact of a spoofing attack on the vehicle’s dynamics and motion. Therefore, predicting the dynamic state of the vehicle becomes crucial. To achieve this, a simple 2 Degrees of Freedom (2 DOF) dynamic bicycle model proposed by Rajamani et al. [29] is employed. Fig. 2 depicts the dynamic bicycle model of a vehicle in a 2-dimensional inertial frame.
@@
Dynamic Bicycle Model
@@
Localization/ Navigation
The GPS enables the vehicle to determine its position and localize to the destination. Since this paper focuses on presenting a GPS intrusion detection system, only GPS-based navigation is considered, and camera or vision sensor-based perception is not taken into account. To ensure safe GPS-guided localization for autonomous vehicles, it is necessary to continuously monitor various parameters such as GPS latitude and longitude, signal quality, GPS Dilution of Precision (DOP), and the number of satellites the vehicle is locked with. The normal behavior of the GPS signal for vehicular localization can be modeled by constantly monitoring the following parameters:
@@
State Estimation
One of the well-established solutions for estimating the state of nonlinear systems is the Extended Kalman Filter or EKF. It integrates measurements from multiple sensors with a system model to estimate the state of the system with improved accuracy [53]. In the context of a GPS-guided AV, the EKF utilizes multiple sensor measurements in conjunction with the dynamic vehicle model to make accurate estimation about the vehicle’s state, including its position, orientation, velocity, etc. This estimation procedure takes into account the nonlinear relationship between the measurements and the position, while also considering measurement noise. EKF state estimator with GPS/INS fusion is capable of detecting sensor anomalies and attacks against autonomous vehicles by monitoring the measurement deviations of multi-sensor readings [48, 55]. A similar GPS/INS fusion using EKF dynamic state estimator has been employed in our vehicle model and experimental testbed. In Section V, it is shown that the GPS-IDS approach is able to detect a GPS spoofing attack faster than the employed EKF-based GPS/INS detection approach, which is supported by experimental validation.
@@
$\textbf{x}_{k+1}=f_{cd}(\textbf{x}_{k},u_{k},W_{k})$ (16)
$\textbf{y}_{k+1}=g_{cd}(\textbf{x}_{k},E_{k})$ (17)
@@
$F=\left.\frac{\partial f_{cd}}{\partial\textbf{x}}\right|_{\hat{x}_{k-1},u_{k}}$ (18)
$G=\left.\frac{\partial g_{cd}}{\partial\textbf{x}}\right|_{\hat{x}_{k}}$ (19)
Thus, the nonlinear system can be transformed into a linear system with an update process by the following equations:
$\textbf{x}_{k+1}=F\textbf{x}_{k}+W_{k}$ (20)
$\textbf{y}_{k+1}=G\textbf{x}_{k}+E_{k}$ (21)
$\textbf{x}_{\left.{k+1}\right|k}=F\hat{\textbf{x}}_{\left.k\right|k}+W_{k}$ (22)
$\textbf{y}_{\left.{k+1}\right|k}=G\hat{\textbf{x}}_{\left.k\right|k}+E_{k}$ (23)
@@
$\hat{\textbf{x}}_{\left.{k+1}\right|{k+1}}=x_{\left.{k+1}\right|k}+\textbf{K}_{\left.{k+1}\right|k}(\textbf{y}_{k+1}-\textbf{y}_{\left.{k+1}\right|k})$ (24)
Motion Planning
@@
From 25, we can say that when the vehicle is traveling along the desired path, $\psi(t)=\psi_{t}(t)$, and $e(t)=0$. When the vehicle deviates from the desired path, $\psi(t)$ differs from $\psi_{t}(t)$ and the $e(t)$ increases accordingly. In general, an increase in the difference between the target yaw angle and the current yaw angle leads to a corresponding amplification in the cross-track error.
Controller
@@
Derived from the discussed software elements, the architecture of the Autonomous Vehicle Behavior Model is depicted in Fig. 4. From the figure, we can observe that the model is established on 14 features— Localization/ Navigation block: Latitude ($lat$), Longitude ($lon$), Horizontal DOP ($hdop$), Vertical DOP ($vdop$), Satellite Lock ($Sat_{lock}$), Satellite Count ($Sat_{count}$); Motion Planning block: Cross-Track Error ($e$); Controller block: Steering Angle ($\delta$); and Vehicle Dynamics block: Longitudinal Position ($x$), Lateral Position ($y$), Yaw Angle ($\psi$), Longitudinal Velocity ($v_{x}$), Lateral Velocity ($v_{y}$), and Yaw Rate ($r$). Data from these features are acquired by the Data-centric Component, and machine learning techniques are employed to model AV behavior and detect anomalies.
III-B Anomaly Behavior Analysis (ABA) Module
Satam et al. have presented an IDS for wireless networks based on Anomaly Behavior Analysis (ABA) [13, 27], which forms the basis of this module. However, we improved it by integrating both physics and data, rather than relying solely on data to detect GPS spoofing on AVs.
[…]
$D(s)=\begin{cases}Abnormal\hskip 28.45274ptif\quad f_{norm}(s,M)>\mathbb{T}\\ Normal\hskip 36.98866ptotherwise\end{cases}$ (28)
Detection occurs when detector $D$ identifies a sample as abnormal, whether it is truly an anomaly or a normal sample misclassified as one. The detection errors are defined over a test set $U_{t}^{R}$ which is a subset of $U^{R}$, $U_{t}^{R}\subseteq U^{R}$. The detector considers two kinds of errors: False Positives and False Negatives. A False Positive detection occurs when a normal sample $s\in N^{R}$ is detected as an abnormal event and is defined as $\varepsilon^{+}=\{s\in N^{R}|D(s)=abnormal\}$, while a False Negative detection occurs when the detector classifies an abnormal sample $s\in A^{R}$ as a normal event (undetected anomalies), that is $\varepsilon^{-}=\{s\in A^{R}|D(s)=normal\}$. The objective of GPS-IDS is to tune the predefined threshold $\mathbb{T}$, so that the overall error is minimized. Particularly, we prioritize the minimization of False Negative errors, as these undetected attacks pose greater risks compared to false alarms in the context of AVs. In the GPS-IDS approach, ABA is applied using machine learning (Experiment 3: V-D3) and the threshold $\mathbb{T}$ is tuned from optimization (Experiment 4: V-D4).